Circular No. 6476

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 Q1 (TABUR) R. Schulz, Max-Planck-Institut für Aeronomie; and G. P. Tozzi, Os-

servatorio Astrofisico di Arcetri, report on observations of this comet with the MPI/ESO 2.2-m telescope at the European Southern Observatory on Sept. 12 and 13: “We found the comet quite peculiar on Sept. 12, with a very high gas/dust ratio. On a low-resolution spectrogram, the cometary emission lines were very intense and there was almost no continuum. Moreover, narrowband CN images indicate strong jet-like features in the coma. Spectra and images taken on Sept. 13 indicate that the dust emission was much greater than on the previous day; by measuring the relative intensity of C_{2} and the adjacent continuum, we found that the gas/dust ratio had decreased by a factor of three in one day, and the overall intensity increased by a factor of about two.”

Visual m_{1} estimates (B = binoculars): Sept. 7.99 UT, 7.9 (A. Baransky, Kiev, Ukraine, 0.11-m reflector); 10.39, 7.8 (R. Keen, Mt. Thorodin, CO, 7×50 B); 15.82, 6.8 (J. Kobayashi, Kumamoto, Japan, 7×50 B); 18.11, 6.3 (R. J. Bouma, Groningen, The Netherlands, 15×80 B; tail 0^{o}.9 long in p.a. 279^{o}); 20.11, 6.3 (K. Hornoch, Lelekovice, Czech Republic, 10×80 B).

COMET C/1996 B2 (HYAKUTAKE) C. B. Cosmovici, Institute for Space Physics, CNR, Frascati; S. Mon-

tebugnoli and A. Orfei, Institute for Radioastronomy, CNR, Medicina; S. Pogrebenko, JIVE, Dwingeloo; and S. Cortiglioni, TESRE, CNR, Bologna, report: “We detected the 22-GHz water-emission line from C/1996 B2 near the time of perihelion, using the 32-m dish of the Medicina Radiotelescope and a programmable high-resolution spectrum analyzer (9-MHz input bandwidth, 8192 channels) as back-end; the field-of-view was 2′. The results are peculiar, with the line being split into two clear and strong components (maximum antenna temperature on May 3 was 130 ± 20 mK), and positively shifted from the zero-velocity point as follows: Apr. 26, 20 and 32 km/s; May 1, 27 and 35; 3, 34 and 41; 4, 38 and 45. A complete analysis of the data exclude contamination by other possible spectral lines and local oscillator position errors. As far as we know, similar behavior of emission lines in comets has never before been reported. In order to understand if this behavior is due to jetting and/or bowshock physics, we encourage other observers to communicate their ground- or space-based observations from

the same period to us at cosmo@ifsi.fra.cnr.it.”

1996 September 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green